OpenStudy (anonymous):

Really need some help understanding this problem?!?! The temperature of a cup of coffee varies according to Newton's Law of Cooling: dT/dt = -k (T-A), where T is the water temperature, A is the room temperature, and k is a positive constant. If the coffee cools from 180°F to 100°F in 10 minutes at a room temperature of 75°F, how long (to the nearest minute) will it take the water to cool to 80°F?

OpenStudy (empty):

Ahhh this is a fun one. Any ideas, how far can you get on your own just doing anything? What have you tried?

OpenStudy (anonymous):

Well in know that it has something to do with the separation of variables, which is just a simple way of solving for a basic differential equation. But the issue lies in interpreting the numbers given and replacing them into the equation to make it solvable. in short, I have nothing

OpenStudy (empty):

Well I think we can do this with an integrating factor or possibly some other way too, this is just an ODE not a PDE so it should be easier to solve. Since we're rounding it's even possible to just estimate some things such as replacing \(\frac{dT}{dt}=\frac{ \Delta T}{\Delta t}\)

OpenStudy (empty):

Ok so here's what we do, substitute in: \[u=T-A\] differentiate to get: \[\frac{du}{dt} = \frac{dT}{dt}\] this just simplifies the differential equation to become: \[\frac{du}{dt} = -ku\] which is separable: \[\int \frac{du}{u}=\int -kdt\] Now you can solve this for T(t), so do this now and I'll help you out. :)

OpenStudy (alekos):

that link is no good

OpenStudy (alekos):

do you know how to solve the integrals?

OpenStudy (anonymous):

yes i do

OpenStudy (alekos):

OK, so whats the next step?

OpenStudy (anonymous):

this is what i'v got so far the left side is likely wrong

OpenStudy (anonymous):

i mean right

OpenStudy (alekos):

Thats all really good! You've found k = 0.14 and C = 105 so now you have the full equation and you can use this to find t when the temperature drops to 75 deg

OpenStudy (alekos):

\[T = 105e ^{-0.14t} + 75\]

OpenStudy (alekos):

I meant find t when the temp drops to 80. So substitute T=80

OpenStudy (anonymous):

did that, was wrong. by the way T = 80 as final temp so it would be \[80 = 105e^(-.14t)+75\]

OpenStudy (alekos):

yep. thats it. now solve for t